# OVTR: End-to-End Open-Vocabulary Multiple Object Tracking with Transformer

### v2

@@ 1 Introduction
Based on the iterative tracking mechanism, we propagate category information across multiple frames, which we refer to as the category information propagation (CIP) strategy. By converting predicted category information into priors for subsequent tracking, it establishes a higher-level flow that enables the continuous tracking of corresponding targets. Additionally, due to the introduction of attention mechanism, potential interference from differing information must also be considered. To address this, we designed two attention protection strategies. These strategies construct masks based on category prediction distributions and the arrangement of queries, effectively intervening to ensure harmonious cooperation between classification and tracking.
@@ 3.2 Leveraging Aligned Queries for Search in Cross-Attention
Dual-Branch Structure. As shown in Fig. 3, in the dual-branch structure, the OFA branch consists of a feed-forward network followed by a box head and an alignment head, while the category text interaction (CTI) branch comprises text cross-attention followed by a feedforward network.
To achieve zero-shot capabilities, we utilize the OFA branch for alignment, distilling visual generalization from the CLIP image encoder and guiding the queries produced by the image cross-attention layer, referred to as aligned queries, to align with CLIP image embeddings. Since the CLIP image embeddings are aligned with the CLIP text embeddings and the text features $E_{\text{txt}}$ originate from the CLIP text embeddings, aligned queries that correspond to certain categories are able to align implicitly with the $E_{\text{txt}}$ of the same categories. This alignment enables the aligned queries to effectively focus on the category information conveyed in $E_{\text{txt}}$ during the text cross-attention in the CTI branch. Consequently, when images and text inputs containing novel categories are provided, the generalized network produces aligned queries that implicitly capture the features of novel categories. This enables the aligned queries to interact effectively with the $E_{\text{txt}}$ of novel categories, ultimately enhancing open-vocabulary perception. The box head is incorporated into the OFA branch to ensure that the features of aligned queries are instance-level.
@@ 3.3 Attention Isolation for Decoder Protection
Category Isolation Strategy. The output features of the CTI branch $O_{\text{txt}}$, undergo dot products with the text features, followed by a softmax operation to produce the category score matrix ${\bm{S}}\in\displaystyle\mathbb{R}^{N\times M}$, where $N$ denotes the number of queries $Q$, and $M$ represents the number of selected categories. We calculate the KL (Kullback-Leibler) divergence (Kullback, 1951) between the category score distributions (vectors in ${\bm{S}}$) of each two predictions in $O_{\text{txt}}$ to form a matrix, called the difference matrix ${\bm{D}}\in\displaystyle\mathbb{R}^{N\times N}$. The specific formula is as follows:
@@
When ${I}_{i,j}$ is ”True”, it indicates that the category information carried by these two queries is substantially different. During the self-attention process, this difference may lead to interference between the queries. Notably, interactions among track queries will be maintained to ensure that the attention mechanism prevents conflicts within predictions corresponding to track queries.
@@ 3.4 Tracking with Category Information Propagation Across Frames
where $P_{*}^{\prime\,t}$ and $C_{*}^{t}$ respectively represent the updated position parts and the content parts of $Q^{\prime\,t}_{*}$, MHA and FFN denote multi-head attention and feed-forward network, and $O_{\text{sa}}$ and $O_{\text{r}}$ are the intermediate attention outputs and residual outputs. We use the sum of $O_{\text{img}}^{*\,t}$ and $P_{*}^{\prime\,t}$ as the queries and keys, while using $O_{\text{img}}^{*\,t}$ alone as the values for the MHA.
This network aggregates the category information from $O_{\text{img}}^{*\,t}$ with historical content, providing category priors in $C_{\text{tr}}^{t+1}$ for the next frame predictions. In this way, category information is propagated to the next frame, enabling multi-frame propagation during the iteration of track queries. Meanwhile, the bounding boxes $B^{t}_{*}$ corresponding to $Q^{\prime\,t}_{*}$ are transformed via sin-cos positional encoding into $P_{\text{tr}}^{t+1}$ of $Q_{\text{tr}}^{t+1}$. We exclusively use the output representations from the OFA branch, instead of the CTI branch, as the OFA branch contains less direct textual information. This can reduce the content gap between detect and track queries. Additionally, since $O_{\text{img}}^{*\,t}$ is derived from the image cross-attention layer, it may more readily align with similar information within that layer when re-entering, thereby facilitating the continuous capture of the target.
@@ 3.5 Optimization
where $\widehat{y}_{\text{tr}}^{i}\mid_{Q_{\text{tr}}}$, $y_{\text{tr}}^{i}$, $\widehat{y}_{\text{det}}^{i}\mid_{Q_{\text{det}}}$, $y_{\text{det}}^{i}$ denote the association predictions, association labels, detection predictions, and non-association labels, respectively. $T_{i}$ is the total number of the targets in $f_{t=i}$.
@@ B.3 Inference Speed Evaluation
OVTR 3.4 31.4 54.4 34.5 5.4 36.6 52.2 37.6 20.1
OVTR-Lite 12.4 30.1 52.7 34.4 3.1 35.6 51.3 37.0 18.6

### v3

@@ 1 Introduction
Based on the iterative tracking mechanism, we propagate category information across multiple frames, which we refer to as the category information propagation (CIP) strategy. By converting predicted category information into priors for subsequent tracking, it establishes a higher-level flow that enables the continuous tracking of corresponding targets. Additionally, due to the introduction of attention mechanism, potential interference among differing information must also be considered. To address this, we designed two attention protection strategies. These strategies construct masks based on category prediction distributions and the arrangement of queries, effectively intervening to ensure harmonious cooperation between classification and tracking.
@@ 3.2 Leveraging Aligned Queries for Search in Cross-Attention
Dual-Branch Structure. As shown in Fig. 3, in the dual-branch structure, the OFA branch consists of a feedforward network followed by a box head and an alignment head, while the category text interaction (CTI) branch comprises text cross-attention followed by a feedforward network.
To achieve zero-shot capabilities, we utilize the OFA branch for alignment, distilling visual generalization from the CLIP image encoder and guiding the queries produced by the image cross-attention layer, referred to as aligned queries, to align with CLIP image embeddings. Since the CLIP image embeddings are aligned with the CLIP text embeddings and the text features $E_{\text{txt}}$ originate from the CLIP text embeddings, aligned queries that correspond to certain categories are able to align implicitly with the $E_{\text{txt}}$ of the same categories. This alignment enables the aligned queries to effectively focus on the category information conveyed in $E_{\text{txt}}$ during the text cross-attention in the CTI branch. Consequently, when images and text inputs containing novel categories are provided, the generalized network produces aligned queries that implicitly capture the features of novel categories. This enables the aligned queries to interact effectively with the $E_{\text{txt}}$ of novel categories, yielding $O_{\text{txt}}$ rich in novel category information to enhance open-vocabulary classification. The box head is incorporated into the OFA branch to ensure that the features of aligned queries are instance-level.
@@ 3.3 Attention Isolation for Decoder Protection
Category Isolation Strategy. The output features of the CTI branch $O_{\text{txt}}$, undergo dot products with the text features $E_{\text{txt}}$, followed by a softmax operation to produce the category score matrix ${\bm{S}}\in\displaystyle\mathbb{R}^{N\times M}$, where $N$ denotes the number of queries $Q$, and $M$ represents the number of selected categories. We calculate the KL (Kullback-Leibler) divergence (Kullback, 1951) between the category score distributions (vectors in ${\bm{S}}$) of each two predictions in $O_{\text{txt}}$ to form a matrix, called the difference matrix ${\bm{D}}\in\displaystyle\mathbb{R}^{N\times N}$. The specific formula is as follows:
@@
When ${I}_{i,j}$ is ”True”, it means the category information carried by these two queries is substantially different. This difference may cause interference between the queries during self-attention. In addition, interactions among $Q_{\text{tr}}$, which have inherited robust category information, will be maintained to ensure that the self-attention mechanism inhibits duplicate predictions derived from $Q_{\text{tr}}$.
@@ 3.4 Tracking with Category Information Propagation Across Frames
where $P_{*}^{\prime\,t}$ and $C_{*}^{t}$ respectively represent the updated position parts and the content parts of $Q^{\prime\,t}_{*}$, MHA and FFN denote multi-head attention and feedforward network, and $O_{\text{sa}}$ and $O_{\text{r}}$ are the intermediate attention outputs and residual outputs. We use the sum of $O_{\text{img}}^{*\,t}$ and $P_{*}^{\prime\,t}$ as the queries and keys, while using $O_{\text{img}}^{*\,t}$ alone as the values for the MHA.
This network aggregates the category information from $O_{\text{img}}^{*\,t}$ with historical content $C_{*}^{t}$, providing category priors in $C_{\text{tr}}^{t+1}$ for the next frame predictions. In this way, category information is propagated to the next frame, enabling multi-frame propagation during the iterations of track queries. Meanwhile, the bounding boxes $B^{t}_{*}$ corresponding to $Q^{\prime\,t}_{*}$ are transformed via sin-cos positional encoding into $P_{\text{tr}}^{t+1}$ of $Q_{\text{tr}}^{t+1}$. We exclusively use the output representations from the OFA branch, instead of the CTI branch, as the OFA branch contains less direct textual information. This can reduce the content gap between detect and track queries. Additionally, since $O_{\text{img}}^{*\,t}$ is derived from the image cross-attention layer, it may more readily align with similar information within that layer when re-entering, thereby facilitating the continuous capture of the target.
@@ 3.5 Optimization
where $\widehat{y}_{\text{tr}}^{i}\mid_{Q_{\text{tr}}}$, $y_{\text{tr}}^{i}$, $\widehat{y}_{\text{det}}^{i}\mid_{Q_{\text{det}}}$, $y_{\text{det}}^{i}$ denote the association predictions, association labels, detection predictions, and unassociated labels, respectively. $T_{i}$ is the total number of the targets in $f_{t=i}$.
@@ B.3 Inference Speed Evaluation
OVTR (Ours) 3.4 31.4 54.4 34.5 5.4 36.6 52.2 37.6 20.1
OVTR-Lite (Ours) 12.4 30.1 52.7 34.4 3.1 35.6 51.3 37.0 18.6